# Large Model Strategic Thinking, Small Model Efficiency: Transferring Theory of Mind in Large Language Models

### v1

@@ Abstract
As the performance of larger, newer Large Language Models continues to improve for strategic Theory of Mind (ToM) tasks, the demand for these state of the art models increases commensurately. However, their deployment is costly both in terms of processing power and time. In this paper, we investigate the feasibility of creating smaller, simulation-ready agents by way of fine-tuning. To do this, we present a large pre-trained model with 20 unique scenarios that combine a social context with a social dilemma, recording its answers, and using them for Q&A fine-tuning on a smaller model of the same family. Our focus is on in-context game-theoretic decision-making, the same domain within which human interaction occurs and that requires both a theory of mind (or a semblance thereof) and an understanding of social dynamics. We find that the fine-tuned smaller language model exhibited significant performance closer to that of its larger relative, and that their improvements extended in areas and contexts beyond the ones provided in the training examples. On average for all games, through fine-tuning, the smaller model showed a %46 improvement in aligning with the behavior of the larger model, with %100 representing complete alignment. This suggests that our pipeline represents an efficient method to transmit some form of theory of mind to smaller models, creating improved and cheaply deployable algorithms in the process. Despite their simplicity and their associated shortcomings and limitations, our findings represent a stepping stone in the pursuit and training of specialized models for strategic and social decision making.
@@ 3 Methods
When testing the performance of pre-trained and fine-tuned model on within-sample scenarios, we run 300 initializations and record the response given by the LLM. Prompts are kept constant across LLMs. In case of LLaMa2-70b, we explicitly ask the algorithm not only to give us its choice of action, but also to include a motivation for it. Although the motivations provided by the LLM might include imprecise reasoning and are unlikely to reflect its actual decision making process, we still chose to include them in the dataset. Given the existing evidence that its final choices should be on average "correct", the accompanying elaborations should reflect and corroborate that with an explanation grounded in natural language. The resulting triplets of system prompt, game prompt and AI response are then collated within a dataset which is used to fine-tune the smaller LLaMa2-7b model. Additionally, as a robustness check, we ask the larger model to output only its choice of action (without motivation) in order to test for stability of preferences. Once the fine-tuning has been performed, we query the fine-tuned model both on the scenarios it has been trained on, and on the new out-of-sample ones using 300 initializations per scenario. For
following the procedure described thus far, sans requiring the larger model to provide us a motivation.

### v2

@@ Abstract
As the performance of larger, newer Large Language Models continues to improve for strategic Theory of Mind (ToM) tasks, the demand for these state of the art models increases commensurately. However, their deployment is costly both in terms of processing power and time. In this paper, we investigate the feasibility of creating smaller, highly-performing specialized algorithms by way of fine-tuning. To do this, we present a large pre-trained model with 20 unique scenarios that combine a social context with a social dilemma, recording its answers, and using them for Q&A fine-tuning on a smaller model of the same family. Our focus is on in-context game-theoretic decision-making, the same domain within which human interaction occurs and that requires both a theory of mind (or a semblance thereof) and an understanding of social dynamics. We find that the fine-tuned smaller language model exhibited significant performance closer to that of its larger relative, and that their improvements extended in areas and contexts beyond the ones provided in the training examples. On average for all games, through fine-tuning, the smaller model showed a 46% improvement in aligning with the behavior of the larger model, with 100% representing complete alignment. This suggests that our pipeline represents an efficient method to transmit some form of theory of mind to smaller models, creating improved and cheaply deployable algorithms in the process. Despite their simplicity and their associated shortcomings and limitations, our findings represent a stepping stone in the pursuit and training of specialized models for strategic and social decision making.
@@ 3 Methods
When testing the performance of pre-trained and fine-tuned models on within-sample scenarios, we run 300 initializations and record the response given by the LLM. Prompts are kept constant across LLMs. For LLaMa2-70b, we explicitly ask the algorithm not only to give us its choice of action, but also to include a motivation for it. Although the motivations provided by the LLM might include imprecise reasoning and are unlikely to reflect its actual decision making process, we still chose to include them in the dataset. Given the existing evidence that its final choices should be on average "correct", the accompanying elaborations should reflect and corroborate that with an explanation grounded in natural language. The resulting triplets of system prompt, game prompt and AI response are then collated within a dataset which is used to fine-tune the smaller LLaMa2-7b model. Additionally, as a robustness check, we ask the larger model to output only its choice of action (without motivation) in order to test for stability of preferences. Once the fine-tuning has been performed, we query the fine-tuned model both on the scenarios it has been trained on, and on the new out-of-sample ones using 300 initializations per scenario.